OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

Write equation with variables J and T. J is 10 years older than T => J = T + 10 J was twice T's age 3 years ago => J-3 = 2(T-3) => J = 2(T-3)+3 Substitute in original equation and find T.

OpenStudy (anonymous):

Huh

OpenStudy (anonymous):

What part of it is confusing?

OpenStudy (anonymous):

What do you mean by substitute in original equation?

OpenStudy (anonymous):

J = T + 10 J = 2(T-3)+3 => 2(T-3)+3 = T + 10 => ......... You can now find T.

OpenStudy (anonymous):

Tim is now 13 year old. and Coach is 23 year. 3 year ago : Tim was 10 year old and Coach was 20 year old. Method of GT is correct.